OpenStudy (anonymous):

A plane starting at rest at the south end of a runway undergoes a constant acceleration of 1.6 m/s/s for a distance of 1600m before takeoff. Part 1: What is the time required for takeoff? Part 2: What is the plane’s velocity at takeoff? **How do we solve these ones? @michele_laino? :)

OpenStudy (michele_laino):

please you have to apply this formula: \[S=\frac{ 1 }{ 2 }a*t ^{2}\] where S=1,600 meters and a=1.6 m/sec^2 please solve that equation for t

OpenStudy (anonymous):

okay! so 1600=1/2(1.6)*t^2 1600=0.8 * t^2 2000=t^2 t=44.72 ? :/

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! :) so that's part 1 correct? what about part 2? :/

OpenStudy (michele_laino):

for part 2 you have to apply this formula: \[v(t)=a*t\] where t is 44.72 sec and a=1.6 m/sec^2

OpenStudy (anonymous):

ah okie! so v(t)= 44.72 * 1.6 = 71.55 m/s ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

awesome!! thank you:)

OpenStudy (michele_laino):

thank you!! :)